# MaterialMVP: Illumination-Invariant Material Generation via Multi-view PBR Diffusion

### v1

@@ Abstract
Figure 1: A series of textured meshes with PBR material generated by MaterialMVP, showcasing exceptional texture generation and realistic physical behavior under environmental illumination. More results can be found on the project page https://zebinhe.github.io/MaterialMVP/.
footnotetext: $\dagger$ Project leader. $\ddagger$ Corresponding author.
@@ 1 Introduction
Generating textures for 3D models is a crucial task in creating visually compelling and realistic digital assets, with numerous studies dedicated to advancing this field [62, 36, 3, 4, 22, 43, 5, 64, 67, 1, 33, 72, 53]. While traditional RGB textures focus on color representation, Physically Based Rendering (PBR) textures take a step further by simulating real-world material properties such as reflectivity, roughness, and light interaction, resulting in more immersive and physically plausible visual experiences.
@@ 2 Related Work
Texture Generation. The development of 3D generation has significantly advanced the creation of high-quality textures for 3D meshes [62, 36, 3, 4, 22, 43, 5, 64, 67, 1, 33, 72, 53, 63]. Extending text-to-image diffusion models[14, 42, 37, 34, 25] provides a foundation for generating detailed and semantically coherent textures from textual descriptions [36, 3, 4, 22, 43, 5, 30, 66, 1, 33]. Building upon this, using image prompts as input enables textures to closely align with specific visual references [64, 67, 72, 68, 63, 20]. Advancing further, 3D priors, such as depth maps [22, 43, 67, 53], normal maps [1, 33, 72], and position maps [1, 72, 64], ensures that the textures are not only visually realistic but also geometrically consistent with the underlying 3D structure.
@@
Multi-view Generation. Multi-view generation [48, 28, 27, 17, 12, 44, 49, 50, 32, 57, 40, 41, 24] has been adopted in texture generation [30, 66, 1, 33, 72, 11, 68] to address issues such as blurring or artifacts. Some methods integrate viewpoint information or 3D priors [8, 19, 29, 56, 58, 15]. Other approaches utilize epipolar geometry to improve consistency between views or reduce memory overhead [24, 18, 21], while another technique tiles multiple view images into a single layer, treating them as a unified input during the diffusion model’s denoising process [40, 41, 51, 23].
@@ 3 Method
As shown in Fig. 2, our MaterialMVP generates PBR textures conditioned on a 3D mesh and an image prompt. The generated albedo map is expected to be free from lighting information, while the metallic and roughness maps should be both accurate and precisely aligned. In this section, we first review the multi-view image generation diffusion model and the principles of PBR-based material modeling, which form the foundation of our approach. Then, we introduce the Consistency-Regularized Training method, designed to disentangle residual lighting information from the albedo map while improving robustness against input variations. Finally, we present our dual-channel material generation framework, which ensures the precise alignment and fidelity of the generated texture maps.
@@ 3.2.1 Reference Pair Selection
For each object, we construct a candidate set $\mathcal{I}$ of 312 images rendered in various camera poses and lighting conditions at elevation angles $\theta_{c}\in\left\{-20^{\circ},0^{\circ},20^{\circ},\text{ random}\right\}$, with 24 azimuth angles $\phi_{c}$ uniformly sampled per elevation. For $\theta_{c}\in\left\{-20^{\circ},0^{\circ},20^{\circ}\right\}$, each pose is rendered under three HDR lighting conditions; for $\theta_{c}=\text{random}$, each pose is rendered under a single point light with random light elevation $\theta_{l}\in\left\{-30^{\circ},0^{\circ},30^{\circ}\right\}$ and light azimuth $\phi_{l}\in\left\{-45^{\circ},0^{\circ},45^{\circ}\right\}$. As shown in Fig. 3, to select a reference pair $(I_{1},I_{2})$, we first choose $I_{1}$ from $\mathcal{I}$ with probability $p$ favoring point light images (in practice, we set $p=0.4$), and then select $I_{2}$ from the subset $\mathcal{S}(I_{1})$ defined as

### v2

@@ Abstract
Figure 1: A series of textured meshes with PBR material generated by MaterialMVP, showcasing exceptional texture generation and realistic physical behavior under environmental illumination. More results and code can be found on https://github.com/ZebinHe/MaterialMVP.
footnotetext: $\dagger$ Project leader.footnotetext: $\ddagger$ Corresponding author.
@@ 1 Introduction
Generating textures for 3D models is a crucial task in creating visually compelling and realistic digital assets, with numerous studies dedicated to advancing this field [66, 40, 3, 4, 23, 47, 5, 68, 71, 1, 37, 76, 57, 20, 21]. While traditional RGB textures focus on color representation, Physically Based Rendering (PBR) textures take a step further by simulating real-world material properties such as reflectivity, roughness, and light interaction, resulting in more immersive and physically plausible visual experiences.
@@ 2 Related Work
Texture Generation. The development of 3D generation has significantly advanced the creation of high-quality textures for 3D meshes [66, 40, 3, 4, 23, 47, 5, 68, 71, 1, 37, 76, 57, 67]. Extending text-to-image diffusion models[14, 46, 41, 38, 29] provides a foundation for generating detailed and semantically coherent textures from textual descriptions [40, 3, 4, 23, 47, 5, 34, 70, 1, 37]. Building upon this, using image prompts as input enables textures to closely align with specific visual references [68, 71, 76, 72, 67]. Advancing further, 3D priors, such as depth maps [23, 47, 71, 57], normal maps [1, 37, 76], and position maps [1, 76, 68], ensures that the textures are not only visually realistic but also geometrically consistent with the underlying 3D structure.
@@
Multi-view Generation. Multi-view generation [52, 32, 31, 17, 12, 48, 53, 54, 36, 61, 44, 45, 26, 25, 28, 27] has been adopted in texture generation [34, 70, 1, 37, 76, 11, 72] to address issues such as blurring or artifacts. Some methods integrate viewpoint information or 3D priors [8, 19, 33, 60, 62, 15]. Other approaches utilize epipolar geometry to improve consistency between views or reduce memory overhead [26, 18, 22], while another technique tiles multiple view images into a single layer, treating them as a unified input during the diffusion model’s denoising process [44, 45, 55, 24].
@@ 3 Method
As shown in Fig. 2, our MaterialMVP generates PBR textures conditioned on a 3D mesh and an image prompt. The 3D mesh is input in the form of normal maps and position maps, which are encoded into latent space and concatenated channel-wise with noise latent as U-Net input. The generated albedo map is expected to be free from lighting information, while the metallic and roughness maps should be both accurate and precisely aligned. In this section, we first review the multi-view image generation diffusion model and the principles of PBR-based material modeling, which form the foundation of our approach. Then, we introduce the Consistency-Regularized Training method, designed to disentangle residual lighting information from the albedo map while improving robustness against input variations. Finally, we present our dual-channel material generation framework, which ensures the precise alignment and fidelity of the generated texture maps.
@@ 3.2.1 Reference Pair Selection
For each object, we construct a candidate set $\mathcal{I}$ of 312 images rendered in various camera poses and lighting conditions at elevation angles $\theta_{c}\in\left\{-20^{\circ},0^{\circ},20^{\circ},\text{ random}\right\}$, with 24 azimuth angles $\phi_{c}$ uniformly sampled per elevation. For $\theta_{c}\in\left\{-20^{\circ},0^{\circ},20^{\circ}\right\}$, each pose is rendered under three random HDR lighting conditions; for $\theta_{c}=\text{random}$, each pose is rendered under a single point light with random light elevation $\theta_{l}\in\left\{-30^{\circ},0^{\circ},30^{\circ}\right\}$ and light azimuth $\phi_{l}\in\left\{-45^{\circ},0^{\circ},45^{\circ}\right\}$. As shown in Fig. 3, to select a reference pair $(I_{1},I_{2})$, we first choose $I_{1}$ from $\mathcal{I}$ with probability $p$ favoring point light images (in practice, we set $p=0.4$), and then select $I_{2}$ from the subset $\mathcal{S}(I_{1})$ defined as
@@ 5 Conclusion
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62372480), in part by the Guangdong Basic and Applied Basic Research Foundation (No. 2023A1515012839), in part by 2025 Tencent AI Lab Rhino-Bird Focused Research Program, in part by Huawei (No. HUAWEI25IS02 and No. 25260020L082), in part by gift fund from Wiener Intelligence (No. HKWITL25IS01), and in part by HKUST-MetaX Joint Lab Fund (No. METAX24EG01-D).